# Learning local equivariant representations for quantum operators

### v3

@@ III.2 Parameterize of Invariant Overlap Operators
where $\phi^{i,l}_{m}$ is the LCAO basis. The overlap matrix also satisfies the equivariance relation defined in Eq.1. Since the equivariant dependency comes fully from the bases, it is possible to rotate them to align with the axis of $\mathbf{r}_{ij}$ to reduce the angular dependence, simplifying the matrix elements further into scalars via the relation [40]:
@@
Here, the dependence changes to $Z_{i}$, $Z_{j}$, and $r_{ij}$ due to the two-center nature of the overlap integrals. Parameterizing these distance-dependent radial functions involves encoding atomic species with radial information and employing a simple MLP to learn the mapping to target scalars.
@@
III.3 The SLEM model
The SLEM model architecture is illustrated in Fig.2. The model maintains a set of features, including hidden features $\mathbf{x}^{ij,L},\mathbf{V}^{ij,L}_{c,l}$, node features $\mathbf{n}^{i,L}_{c,l}$ and edge features $\mathbf{e}^{ij,L}_{c,l}$ of hidden layer ($L$). Specifically, the hidden features consist of a scalar channel $\mathbf{x}^{ij,L}$ and a tensor channel $\mathbf{V}^{ij,L}_{c,l}$. The scalar channel is initialized as an embedded vector containing two-body information, including the atomic species and the radial distance of the atom pair. These initialized two-body radial features are then processed through an MLP to the invariant Slater-Koster-like parameters of overlap. For equivariant quantum operators, such as the Hamiltonian and density matrix, the scalar and tensor channels interact to generate an ordered atom-pair representation, which is used to construct local node and edge representations. After iterative updates, the representation is scaled by the statistical norm and biases, thereby achieving the final prediction.
III.3.1 Feature initialization
Firstly, the initial scalar hidden feature is computed from the two-body embeddings of atomic species $Z_{i}$ and $Z_{j}$, and the radial distance, as follows:
$\mathbf{x}^{ij,L=0}=\textbf{MLP}_{\text{2-body}}\left(\mathbf{1}(Z_{i})||\mathbf{1}(Z_{j})||\textbf{B}(r_{ij})\right)\cdot u(r_{ij})$
Here, atom species are embedded with one-hot vectors denoted as $\mathbf{1}(Z)$, and a set of trainable Bessel bases $\textbf{B}(r_{ij})$ is utilized to encode the distance $r_{ij}$ between atoms $i$ and $j$. $u(r_{ij})$ are envelope functions [41] to add explicit radial dependence. Subsequently, the edge and hidden features are initialized as weighted spherical harmonics of relative edge vectors:
$\mathbf{V}^{ij,L=0}_{c,l}$ $=w_{c,l}\left(L_{N}\left(\mathbf{x}^{ij,L=0}\right)\right)\mathbf{Y}^{ij}_{l}$
@@ III.3.1 Feature initialization
$\mathbf{n}_{c,l}^{i,L=0}=\textbf{Linear}\left(\frac{1}{\sqrt{N_{avg}}}\sum_{j\in\mathcal{N}(i)}\mathbf{e}^{ij,L=0}_{c,l}\right)$
@@
III.3.2 Speed up tensor product
@@
$\begin{split}\mathbf{f}^{ij}_{c_{3},l_{3}}&=\tilde{\mathbf{f}}^{ij}_{c_{1},l_{1}}\otimes w^{ij}_{c_{2},l_{2}}\mathbf{Y}_{l_{2}}^{ij}\\ &=\sum_{c_{1},l_{1},l_{2}}\tilde{w}_{c_{1},l_{1},l_{2}}^{ij}\sum_{m_{1},m_{2}}C_{(l_{1},m_{1})(l_{2},m_{2})}^{(l_{3},m_{3})}\\ &\cdot f^{ij}_{c_{1},l_{1},m_{1}}Y_{l_{2},m_{2}}^{ij}\\ \end{split}$ (4)
Here, $\tilde{w}_{c_{1},l_{1},l_{2}}^{ij}=\sum_{c_{2}}w_{c_{1},c_{2},l_{1},l_{2}}w^{ij}_{c_{2},l_{2}}$ are edge-specific parameters for each tensor product operation. Performing such tensor products on high-order features is computationally intensive. Therefore, we applied the recently developed SO(2) [19] convolution to simplify, reducing the computation and storage complexity from $O(l_{max}^{6})$ to $O(l_{max}^{3})$. The simplification idea is intuitive. $\mathbf{Y}_{l_{2},m_{2}}^{ij}$ are sparse tensors if rotated to align with the edge $ij$, which is nonzero only for $m_{2}=0$. Therefore, it is easier to compute the tensor production in the direction of edge $ij$, and rotate inversely the output afterwards. This step removes the $m_{2}$ index from the summation in Eq. LABEL:tp. Additionally, considering Clebsch-Gordan coefficients with $m_{2}=0$, we find that $C_{(l_{1},m_{1})(l_{2},0)}^{(l_{3},m_{3})}=0$ except for $m_{3}=\pm m_{1}$. This allows further reduction of the summation in Eq. LABEL:tp by replacing ±$\pm m_{1}$ with a single index $m$. Then the operations can be reformulated formally as:
III.3.3 Hidden updates
@@
Unlike most MPNN, the hidden states $\mathbf{x}^{ij}$ and $\mathbf{V}^{ij}_{c,l}$ in SLEM is only a function of the local environment of centre atom $i$, the shift vector of edge $ij$, and the information of atom $j$. Such a design excludes neighbours of $j$ into hidden states.
After the tensor production, the output features $\tilde{\mathbf{V}}^{ij,L}_{c,l}$ will be passed through the gated non-linearity [41], and transformed by an E3linear [42] layer to mix up the information across different channels. The new hidden feature will be multiplied by the weights learned from normalized scalar features to include the radial information explicitly and return as the updated feature $\mathbf{V}^{ij,L}_{c,l}$.
@@
Containing the explicit decaying envelope function $u(r_{ij})$, and the many-body interactions of scalar and tensor features, the sclar hiddien states $\mathbf{x}^{ij,L}$ can be very useful to describing the explicit decaying behaviour of each edge irreps feature along the radial distance.
III.3.4 Node updates
The strict local node representation $\mathbf{n}^{ij,L}_{c,l}$ can be constructed naturally from the many-body interactive tensor features $\mathbf{V}^{ij,L}_{c,l}$. We follows the MPNN style to create the message from node-$j$ to node-$i$. Formally:
@@ III.3.4 Node updates
Here again, we exclude the neighbouring information of atom-$j$ in $\mathbf{m}^{ij,L}_{c_{3},l_{3}}$ via this partial updates of $\mathbf{V}^{ij,L}_{c,l}$. However, all necessary interactions are contained in this format, including the interactions of atom-$j$, bond $i-j$, atom-$i$, and the environment of $i$.
Each message then is passed through a gated activation and E3Linear layer, and weighted separately by weights learnt from the hidden scalar features, by:
$\mathbf{n}^{ij,L}_{c_{3},l_{3}}=\frac{1}{N_{avg}}\sum_{j\in\mathcal{N}(i)}w^{ij}_{c_{3},l_{3}}\mathbf{m}^{ij,L}_{c_{3},l_{3}}$ (8)
The weights here are different from those in hidden updates since it is learnt from $\mathbf{x}^{ij,L}$ directly without passing through a normalization. Therefore, the absolute radial decay is enforced in the weights, giving strong priors that the message from a closer atom is often more important while maintaining flexibility. Meanwhile, combining the updates of hidden scalar $\mathbf{x}^{ij,L}$, we can see the weights $w^{ij}_{c_{3},l_{3}}$ are dependent on the features $\mathbf{V}^{ij,L}$ and $\mathbf{n}^{i,L-1}$, as shown in Eq.6. Such style aligns surprisingly with the equivariant graph attention mechanism [16, 17] that shows powerful expressibility in various tasks. Here ${w}^{ij}_{c_{3},l_{3}}$ corresponding to an attention score computed from $\mathbf{V}^{ij,L}$ and $\mathbf{n}^{i,L-1}$. Therefore, through this update, the node features’ dependencies are kept strictly local.
III.3.5 Edge updates
The locality of edge features $\mathbf{e}^{ij,L}_{c,l}$ can be naturally preserved as long as the node features are localized strictly. By mixing the information of node features on both sides, localized updates can be formulated as:
$\tilde{\mathbf{e}}^{ij,L}_{c_{3},l_{3}}=\left(\mathbf{n}^{i,L-1}||\mathbf{V}^{ij,L}||\mathbf{n}^{j,L-1}\right)_{c_{1},l_{1}}\otimes w_{c_{2},l_{2}}^{ij,L}\mathbf{Y}_{l_{2}}^{ij}$
@@ III.3.5 Edge updates
$\mathbf{e}^{ij,L}_{c_{3},l_{3}}=w^{ij}_{c_{3},l_{3}}\tilde{\mathbf{e}}^{ij,L}_{c_{3},l_{3}}$
Here we reframe the updating rules of SLEM with the MPNN framework, which looks like this:
$\mathbf{e}^{ij,L}$ $=\mathcal{N}_{L}\left(\mathbf{n}^{i,L-1},\mathbf{V}^{ij,L},\mathbf{n}^{j,L-1}\right)$
$\mathbf{m}^{ij,L}$ $=M_{L}\left(\mathbf{n}^{i,L-1},\mathbf{V}^{ij,L}\right)$
Where $\mathcal{V}_{L}$ is the neural network for hidden feature construction. This update scheme constructs many-body interactions to build equivariant edge and node features while preserving the absolute locality by excluding atoms outside the constant cutoff radius. Such a scheme, in principle, would have much better transferability, data-efficiency, as well as scalability. Most importantly, it is possible to parallelize such update computations, allowing for training and inference on multiple devices. This is also vital for inferencing on very large atomic structures. The following section will focus on validating the effectiveness of this framework via learning equivariant DFT Hamiltonians and density matrices and overlap, which have very high-order tensors and need extremely high prediction accuracy. We demonstrate our method is much better than the previously reported ones in training and inference speed, transferability and the potential power to parralize the inference on multi-GPU devices.
@@
Table 2: MAE (in meV) for Hamiltonian matrix predictions using SLEM and DeepH-E3 models [8] on materials with LCAO basis up to $f$ and $g$ orbitals. Numbers in parentheses indicate parameter count. The MAE of HfO2 in DeepH-E3 is absent due to out-of-memory errors.
@@
SLEM charge density model
@@
4.   Results
IV.1 Benchmark the accuracy and data-efficiency
@@
We benchmark our model’s performance in fitting Hamiltonian, density matrix, and overlap matrix using a diverse dataset. For the Hamiltonian, the datasets include systems with up to $d$ orbitals, encompassing 2D systems like single-layer MoS2 and graphene from reported datasets [11], as well as 3D bulk silicon generated in this work. Given SLEM’s proficiency with high-order spherical tensors, we also train it on our generated datasets of bulk GaN and HfO2 systems, which include $f$ and $g$ orbitals, respectively. Structures in the Si, GaN, and HfO2 systems are sampled via molecular dynamics using neural network potentials [34]. For density matrix and overlap matrix evaluations, we test the SLEM model exclusively on the datasets of Si, GaN, and HfO2, as the reported datasets for MoS2 and graphene lack density and overlap matrix data.
@@
Table LABEL:hamiltonian presents a comparison of mean absolute error (MAE) values in Hamiltonian prediction for graphene, MoS2, and Si systems, whose LCAO basis extends up to $d$ orbitals, among the SLEM, DeepH-E3 [8], and HamGNN [10] methods. Our SLEM model achieves state-of-the-art accuracy, exhibiting the lowest MAE across all systems. Notably, this high accuracy is achieved with a relatively small model size of only 0.7 million (M) trainable parameters. Furthermore, we extended our comparison between the SLEM and DeepH-E3 methods to include GaN and HfO2 systems, where the LCAO basis extends up to $f$ and $g$ orbitals, respectively. As shown in Table LABEL:hamiltonian_fg, the SLEM model consistently presents the lowest MAE, further demonstrating its high accuracy and versatility across different orbital complexities. Fig. 3 illustrates the band structure of silicon structures computed directly from the SLEM-predicted Hamiltonian, where the eigenvalues are indistinguishable from those obtained using DFT. For the density matrix, fitting results are presented in Table LABEL:density. The results demonstrate very high accuracy (order of 1e-5), approaching the machine precision limit of float32 numbers. In Fig. 3, we use the trained model to predict the density matrix and visualize its real-space distribution. This capability is particularly important for applications such as charge distribution analysis or tracking electron transfer via methods like differential charge density. Furthermore, the predicted density can be directly used for non-self-consistent field (NSCF) DFT calculations. The resulting band structure for silicon, as an example, is highly accurate and matches the DFT output, with a MAE of only 1.09 meV in eigenvalues compared to self-consistent DFT results. The overlap matrix, represented by invariant parameters in our SLEM model, achieves exceptionally high accuracy as demonstrated in Table LABEL:overlap, approaching the machine precision limit for float32 numbers. Notably, our simplified parameterization enables this high accuracy with only a minimal increase in model complexity. For instance, in a silicon model designed to fit only the Hamiltonian, our typical parameter count is 0.7 M. The inclusion of overlap matrix prediction adds merely 0.01 M parameters, representing an increase of only about 1.4% in total model size.
Table 5: Comparison of validation MAE (in meV) for SLEM model and DeepH-E3 [8] method trained on randomly split datasets [11] with varying training ratios. Model settings align with those in Table LABEL:hamiltonian.
@@
Figure 4: Comparison of time and memory consumption for different tensor-product implementations. (a) Time consumption vs. angular momentum ($l$) for different models, including the SO(2)-based SLEM model (triangles) with and without radial part ($r$), DeepH-E3 (cross) [8], and E3NN (square) [42] models. Inset: Log-scale fit with slopes of 1.2 for the SLEM model and 3.7 for the other two models. (b) Memory consumption vs. $l$. The SLEM model demonstrates over two orders of magnitude improvement in both time and memory efficiency compared to DeepH-E3 and E3NN.
@@
Additionally, the strict localization scheme of our SLEM model confers superior data efficiency, requiring fewer DFT-calculated data points for training. To quantify this efficiency, we conducted an experiment using randomly split subsets of the original training data, comprising 20%, 40%, 60%, and 80% of the full dataset. We trained both the SLEM model and DeepH-E3 method on these subsets and evaluated their performance on consistent validation sets. Table LABEL:data-efficiency presents the results, demonstrating SLEM’s maintained high accuracy across all subsets, thus highlighting its remarkable data efficiency. This data efficiency implies that SLEM users can generate smaller, more cost-effective training sets. Moreover, SLEM’s excellent data efficiency and transferability make it particularly well-suited as a backbone for developing universal DFT models, especially for systems involving heavy elements.
IV.2 Efficiency and Scalability
Many interesting and important properties in material science, chemistry and biology appear with systems composed of heavy atoms. The inclusion of heavy atoms would introduce very high-order spherical tensors when representing quantum operators such as the DFT Hamiltonian. Therefore, scaling to such systems would be quite a challenge because the tensor product used to construct complex spherical tensors scales as $O(L^{6})$. It is very hard to train and inference on systems with heavy atoms using conventional tensor production. Moreover, inferring large material systems while training with small structures is particularly valuable. It therefore requires parallelising the model inference by assigning partitions of the whole atomic structure to multiple GPU workers. However, this is very difficult for most of the currently available models. Since the increasing receptive fields from the iterative updates of the graph would increase the minimal size of each partitioned subgraph, which makes such partition less effective. The SLEM model addresses these challenges by efficiently constructing high-order tensor products and assisting parallelization through its strict locality.
Firstly, for efficiency, the implementation of SO(2) convolution reduces the tensor product computational complexity from $O(L^{6})$ to $O(L^{3})$, which is then further reduced by the parallelization of matrix operations to near $O(L)$ benefiting from PyTorch. In Fig. 4, we compare the wall time and GPU memory consumed by tensor product operations using SO(2) convolution and the normal method employed in DeepH-E3 [8] and E3NN [42]. Clearly, tensor production via SO(2) convolution is far more efficient than conventional methods. This reduction of cost makes our method very efficient, which is capable of handling all possible basis choices in LCAO DFT. We also record the memory and time required when training on some of the typical systems and compare with that of DeepH-E3 [8]. The results are displayed in Fig. 5. We can see, that in all systems, our model trains faster in time while requiring less memory. Notably, the larger the basis set, the greater the advantage our method has in both memory and time efficiency.

### v4

@@ 3.2 Parameterize of Invariant Overlap Operators
where $\phi^{i,l}_{m}$ is the obital from LCAO bases. The overlap matrix also satisfies the equivariance relation. Since the equivariant dependency comes fully from the bases, it is possible to rotate them to align with the axis of $\mathbf{r}_{ij}$, reducing the angular dependence. Therefore, we can further simplify the matrix elements into scalars via the relation Podolskiy & Vogl (2004):
@@
Here, the dependency changes to $Z_{i}$, $Z_{j}$, and $r_{ij}$, indicating the two-center nature of the overlap integrals. Parameterizing these distance-dependent radial functions involves encoding atomic species with radial information and a simple MLP to learn the mapping to target scalars.
@@
3.3 The SLEM model
The SLEM model architecture is illustrated in Fig. 2. The model maintains a set of features, including hidden features $\mathbf{x}^{ij,L},\mathbf{V}^{ij,L}_{c,l}$, node features $\mathbf{n}^{i,L}_{c,l}$ and edge features $\mathbf{e}^{ij,L}_{c,l}$ of hidden layer ($L$). Specifically, the hidden features consist of a scalar channel $\mathbf{x}^{ij,L}$ and a tensor channel $\mathbf{V}^{ij,L}_{c,l}$. The scalar channel is initialized as an embedded vector containing two-body information, including the atomic species and the radial distance of the atom pair. These initialized two-body radial features are then mapped by an MLP to the invariant SK parameters for overlap. For equivariant quantum operators, such as the Hamiltonian and density matrix, the scalar and tensor channels interact to generate an ordered atom-pair representations, which are used to construct local node and edge representations. After iterative updates, the representation is scaled by the statistical norm and biases, thereby achieving the final prediction.
3.3.1 Feature initialization
Firstly, the initial scalar hidden feature is computed from the two-body embeddings of atomic species $Z_{i}$ and $Z_{j}$, and the radial distance $r_{ij}$, as follows:
$\mathbf{x}^{ij,L=0}=\textbf{MLP}_{\text{2-body}}\left(\mathbf{1}(Z_{i})||\mathbf{1}(Z_{j})||\textbf{B}(r_{ij})\right)\cdot u(r_{ij})$ (8)
Here, ||||| | stands for vector concatenation, atom species are embedded with one-hot vectors denoted as $\mathbf{1}(Z)$, and a set of trainable Bessel bases $\textbf{B}(r_{ij})$ is utilized to encode the distance $r_{ij}$ between atoms $i$ and $j$. $u(r_{ij})$ are envelope functions Batzner et al. (2022) to add explicit radial dependence. Subsequently, the edge and hidden features are initialized as weighted spherical harmonics of relative edge vectors:
$\mathbf{V}^{ij,L=0}_{c,l}$ $=w_{c,l}\left(L_{N}\left(\mathbf{x}^{ij,L=0}\right)\right)\mathbf{Y}^{ij}_{l}$ (9)
@@ 3.3.1 Feature initialization
$\mathbf{n}_{c,l}^{i,L=0}=\textbf{Linear}\left(\frac{1}{\sqrt{N_{\text{avg}}}}\sum_{j\in\mathcal{N}(i)}\mathbf{e}^{ij,L=0}_{c,l}\right)$ (10)
@@
3.3.2 Speed up tensor product
@@
$\begin{split}\mathbf{f}^{ij}_{c_{3},l_{3}}&=\tilde{\mathbf{f}}^{ij}_{c_{1},l_{1}}\otimes w^{ij}_{c_{2},l_{2}}\mathbf{Y}_{l_{2}}^{ij}=\sum_{c_{1},l_{1},l_{2}}\tilde{w}_{c_{1},l_{1},l_{2}}^{ij}\sum_{m_{1},m_{2}}C_{(l_{1},m_{1})(l_{2},m_{2})}^{(l_{3},m_{3})}f^{ij}_{c_{1},l_{1},m_{1}}Y_{l_{2},m_{2}}^{ij}\end{split}$ (11)
Here, $\tilde{w}_{c_{1},l_{1},l_{2}}^{ij}=\sum_{c_{2}}w_{c_{1},c_{2},l_{1},l_{2}}w^{ij}_{c_{2},l_{2}}$ are edge-specific parameters for each tensor product operation. Performing such tensor products on high-order features is computationally intensive. Therefore, we applied the recently developed SO(2) Passaro & Zitnick (2023) convolution to reduce the computation and storage complexity from $O(l_{\text{max}}^{6})$ to $O(l_{\text{max}}^{3})$ which we refer to the appendix.
3.3.3 Hidden updates
@@
Unlike most MPNN, the hidden states $\mathbf{x}^{ij}$ and $\mathbf{V}^{ij}_{c,l}$ in SLEM depend only on the local environment of centre atom $i$, the shift vector ${\bm{r}}_{ij}$, and the atomic type and coordinate informations of atom $j$. Such a design excludes neighbours of $j$ into hidden states.
@@
The scalar hidden states $\mathbf{x}^{ij,L}$ incorporate an explicit decaying envelope function $u(r_{ij})$ and many-body interactions of scalar and tensor features. This formulation effectively captures the decay behavior of each edge irrep feature as a function of radial distance.
3.3.4 Node updates
The strict local node representation $\mathbf{n}^{ij,L}_{c,l}$ can be constructed naturally from the many-body interactive tensor features $\mathbf{V}^{ij,L}_{c,l}$. We follow the MPNN style to create the message from node $j$ to node $i$. Formally:
@@ 3.3.4 Node updates
Here again, we exclude the neighbouring information of atom $j$ in $\mathbf{m}^{ij,L}_{c_{3},l_{3}}$ via the partial updates of $\mathbf{V}^{ij,L}_{c,l}$, while maintaining necessary interactions. Each message then is passed through a gated activation and E3Linear layer, weighted separately by weights learnt from the hidden scalar features, and aggregated to update the node feature by:
$\mathbf{n}^{ij,L}_{c_{3},l_{3}}=\alpha\cdot\mathbf{n}^{ij,L-1}_{c_{3},l_{3}}+\frac{\sqrt{1-\alpha^{2}}}{N_{avg}}\sum_{j\in\mathcal{N}(i)}w^{ij}_{c_{3},l_{3}}\mathbf{m}^{ij,L}_{c_{3},l_{3}}$ (15)
Here $\alpha$ ranged from 0 to 1. The weights here differ from those in hidden updates as they are directly learnt from $\mathbf{x}^{ij,L}$ without normalization. Therefore, the absolute radial decay is enforced in the weights, providing a strong prior that the messages from atoms at shorter distances are generally more significant. Meanwhile, the update of $\mathbf{x}^{ij,L}$, and consequently $w^{ij}_{c_{3},l_{3}}$, depends on the features $\mathbf{V}^{ij,L}$ and $\mathbf{n}^{i,L-1}$, as shown in Eq.13. This structure aligns with the equivariant graph attention mechanism Liao & Smidt (2022); Liao et al. (2023) which has demonstrated powerful expressibility in various tasks. Here ${w}^{ij}_{c_{3},l_{3}}$ corresponding to an attention score computed from $\mathbf{V}^{ij,L}$ and $\mathbf{n}^{i,L-1}$. Therefore, through this update, the dependencies of node features are strictly local.
3.3.5 Edge updates
The locality of edge features $\mathbf{e}^{ij,L}_{c,l}$ can be naturally preserved as long as the node features are strictly local. By mixing the information of node features on both sides, localized edge updates can be formulated as follows:
$\tilde{\mathbf{e}}^{ij,L}_{c_{3},l_{3}}=\left(\mathbf{n}^{i,L-1}||\mathbf{V}^{ij,L}||\mathbf{n}^{j,L-1}\right)_{c_{1},l_{1}}\otimes w_{c_{2},l_{2}}^{ij,L}\mathbf{Y}_{l_{2}}^{ij}$ (16)
@@ 3.3.5 Edge updates
$\mathbf{e}^{ij,L}_{c_{3},l_{3}}=\alpha\cdot\mathbf{e}^{ij,L-1}_{c_{3},l_{3}}+\sqrt{1-\alpha^{2}}\cdot w^{ij}_{c_{3},l_{3}}\tilde{\mathbf{e}}^{ij,L}_{c_{3},l_{3}}$ (17)
The following section focuses on validating the effectiveness of this framework via learning equivariant DFT Hamiltonians, density matrices and overlap.
4 Results
4.1 Benchmark the accuracy and data-efficiency
@@
Table 2: MAE (in meV) for Hamiltonian matrix predictions using SLEM and DeepH-E3 models Gong et al. (2023) on materials with LCAO basis up to $f$ and $g$ orbitals. Numbers in parentheses indicate parameter count. The MAE of HfO2 in DeepH-E3 is absent due to out-of-memory errors.
@@
          SLEM density matrix model
@@
We evaluate our model’s performance in fitting Hamiltonian, density matrix, and overlap matrix using diverse datasets. For Hamiltonian, we use systems with up to $d$ orbitals, including 2D systems of monolayer MoS2 and graphene from existing datasets Li et al. (2022), as well as 3D bulk silicon generated for this study. To test SLEM’s capability with high-order tensors, we also train the models on generated datasets of bulk GaN and HfO2 systems, which include $f$ and $g$ orbitals. Structures in the Si, GaN, and HfO2 systems are sampled via molecular dynamics using neural network potentials Wang et al. (2018). For density matrix and overlap matrix evaluations, we focus exclusively on the datasets of Si, GaN, and HfO2 datasets, as the reported datasets for MoS2 and graphene lack density and overlap matrix data. Our model supports an atom-specific $r_{cut}$ setting, each value corresponding to the radial basis cutoff of the numerical atomic orbitals, which aligns with the setting in the compared method.
@@
Table 5: Comparison of validation MAE (in meV) for SLEM model and DeepH-E3 Gong et al. (2023) method trained on randomly split datasets Li et al. (2022) with varying training ratios. Model settings align with those in Table LABEL:hamiltonian.
@@
Table LABEL:hamiltonian presents a comparison of mean absolute error (MAE) values in Hamiltonian prediction for graphene, MoS2, and Si systems, whose LCAO basis extends up to $d$ orbitals, among the SLEM, DeepH-E3 Gong et al. (2023), and HamGNN Zhong et al. (2023) methods. Our SLEM model achieves state-of-the-art accuracy, exhibiting the lowest MAE across all systems. Notably, this high accuracy is achieved with a relatively small model size of only 0.7 million (M) trainable parameters. Furthermore, we extended our comparison between the SLEM and DeepH-E3 methods to include GaN and HfO2 systems, where the LCAO basis extends up to $f$ and $g$ orbitals, respectively. As shown in Table LABEL:hamiltonian_fg, the SLEM model consistently presents the lowest MAE, further demonstrating its high accuracy and versatility across different orbital complexities. Fig. 3 illustrates the band structure of silicon structures computed directly from the SLEM-predicted Hamiltonian, where the eigenvalues are indistinguishable from those obtained using DFT. For the density matrix, fitting results are presented in Table LABEL:density. The results demonstrate very high accuracy (order of 1e-5), approaching the machine precision limit of float32 numbers. In Fig. 3, we use the trained model to predict the density matrix and visualize its real-space distribution. This capability is particularly important for applications such as charge distribution analysis or tracking electron transfer. Furthermore, the predicted density can be directly used for non-self-consistent field (NSCF) DFT calculations. The resulting band structure for silicon, as an example, is highly accurate and matches the DFT output, with a MAE of only 1.09 meV in eigenvalues compared to self-consistent DFT results. The overlap matrix, represented by invariant SK parameters in our SLEM model, achieves exceptionally high accuracy as demonstrated in Table LABEL:overlap, approaching the machine precision limit for float32 numbers. Notably, our simplified parameterization enables this high accuracy with only a minimal increase in model complexity. For instance, in a silicon model designed to fit only the Hamiltonian, our typical parameter count is 0.7 M. The inclusion of overlap matrix prediction adds merely 0.01 M parameters, which is about 1.4% in total model size.
@@
Additionally, the strict localization scheme of our SLEM model confers superior data efficiency, requiring fewer DFT-calculated data points for training. To quantify this efficiency, we conducted an experiment using randomly split subsets of the original data, comprising 20%, 40%, 60%, and 80% of the full training set. We trained both the SLEM model and DeepH-E3 method on these subsets and evaluated their performance on consistent validation sets. Results in table LABEL:data-efficiency demonstrate SLEM’s high accuracy across all subsets, thus highlighting its remarkable data efficiency. This data efficiency implies that SLEM users can generate smaller, more cost-effective training sets. Moreover, SLEM’s excellent data efficiency and transferability make it particularly well-suited as a backbone for developing universal DFT models, especially for systems involving heavy elements.
4.2 Efficiency and Scalability
Figure 4: Comparison of time and memory consumption for different tensor-product implementations. (a) Time consumption vs. angular momentum ($l$) for different models, including the SO(2)-based SLEM model (triangles) with and without radial part ($r$), DeepH-E3 (cross) Gong et al. (2023), and E3NN (square) Geiger & Smidt (2022) models. Inset: Log-scale fit with slopes of 1.2 for the SLEM model and 3.7 for the other two models. (b) Memory consumption vs. $l$. The SLEM model demonstrates over two orders of magnitude improvement in both time and memory efficiency compared to DeepH-E3 and E3NN.